If a game were "fair," the payoff on a bet would be the same as the odds for the event. In one game, the odds for winning are 1:13. If the game were "fair," what would the payoff be for a $5 bet? Of course, game in casinos are designed to make a profit for the casino investors. Supposing the casino makes the payoff at 1:11 odds, what profit does the casino make on your winning bet?
Answer by edjones(8007)

1-13 means that out of every 14 events the bettor loses 13 $1 bets and wins one payoff of $13.
If the payoff is 11 the bettor loses 13 $1 bets and wins 1 payoff of $11 for a net loss of $2.
2/14=1/7= 14 2/7% profit the casino makes.
.
Ed